# A Survey on Monocular Re-Localization: From the Perspective of Scene Map Representation

### v1

@@ 3.1 Visual Place Recognition
Meanwhile, many other attempts are made by scholars to boost the representative ability of global feature. GeM GeM designs a trainable generalized-mean pooling layer that generalizes maximum and average pooling for feature aggregation. TransVPR transvpr could aggregate task-relevant features by self-attention mechanism in Transformers transformer . MixVPR mixvpr incorporates a global relationship between elements in each feature map in cascaded feature mixing modules. AnyLoc keetha2023anyloc leverages an off-the-shelf self-supervised foundation model features (such as DINOv2 oquab2023dinov2 ) with no VPR-specific training or finetuning. Although global feature can compactly represent images, the retrieval time of GF-VPR methods will linearly increase with the size of the database. And the robustness against viewpoint changes and dynamic occlusions of global features is also inferior than local feature DSFeat .
Figure 5: Four typical type of image representation in VPR researches, including a) image global feature, b) image global feature, c) image sequence feature, and d) image semantics.
Local Feature-based VPR (LF-VPR): A local feature algorithm extracts massive key points and corresponding descriptors from images. Then, the extracted local features are utilized to retrieve the best-matched reference frames. Traditional local features are hand-crafted based on human knowledge sift ; orb ; surf . Recently, many DNN-based local features have been adopted in VPR tasks. Early researches directly apply off-the-sheld DNN model pretrained by other computer vision tasks as local feature extraction module BoCNF ; PlaceRW ; convnet_pr . Then, specific DNN-based local features are introduced in LF-VPR researches and greatly improve the performance superpoint .
Directly matching between local features across two images is costly, local features need to be further processed to measure image similarity score. One common way is to aggregate extract local features to a compact global feature, e.g., VLAD vlad ; multivlad , ASMK ASMK , and BoW bow , and retrieve reference frames as GF-VPR methods. As typical solutions. DBoW dbow-iros ; dbow-tro extracts binary local features ORB orb from images and build vocabulary to aggregate local features so that an image can be compactly represented by the occurrences of visual words in the vocabulary (some general local features). Yue et al. yue1 ; yue2 improved it by introducing SuperPoint features superpoint . LRO LRO also utilizes SuperPoint features superpoint but it adopts an ASMK-based retrieval framework ASMK . In dynamic scene, local features located at dynamic objects/regions will confuse VPR methods, so Chen et al. slcd removed inconsistent dynamics by exploiting instance-level semantics. But semantics are hard to accurately judge the actual motion state of the objects, so DSFeat DSFeat learns to select stable and disciminative local features based on not only semantics but also attention information. These LF-VPR methods need to build offline vocabulary/codebook to aggregate local features, limiting the generalization ability. Some methods propose to build vocabularies in an online or incremental manner. In AVBoW , the authors proposed a novel updating strategy to automatically build and update vocabulary so that retrieval can still be performanced as dbow-iros ; dbow-tro .
Different from the solution aggregating local features and retrieving by aggregated feature similarity, some proposals tends to adopt probabilistic scheme in Eqn. 2 for finding matched reference frame. FAB-MAP fabmap and FAB-MAP 2.0 fabmap2 train a Chow-Liu tree to learn the co-visibility of features and then estimate the matching probability of candidates. IBuILD ibuild tracks BRISK features brisk across consecutive frames to incrementally generate vocabulary, and perform VPR by a likelihood function. iBoW-LCD ibow-lcd applies Bayesian framework by measuring the score of local features occurred in current image. LiPo-LCD lipolcd achieves further improvement by adding line-wise local feature in human-made environments. Tsintotas et al. AVWs dynamically segmented the incoming image stream to particular places and assigned places with visual words by using an on-line clustering algorithm . Then in BoTW , Tsintotas et al. improved the former work by tracking SURF features surf by KLT point tracker KLT to obtain robust tracked words and assigned them to the corresponding locations along the map, then VPR is performed by a binomial probability density function.
Figure 6: Three G-RPR solution proposed in arnold2022mapfree , namely, a) first solve scaleless relative pose by 2D-2D matches then recover scale by 3D-3D matches using estimated depth, b) first lift 2D-2D matches to 2D-3D matches by depth then estimate scaled pose by PnP+RANSAC solver, c) lift 2D-2D matches to 3D-3D matches using depth then estimate scaled pose.
@@
Some methods exploit geometrical information from local features as a re-ranked matching score for matching candidates in Eqn. 2. Built upon NetVLAD netvlad , Patch-NetVLAD patch-netvlad re-orders candidates by cross matching of patch-level local DNN features with geometrical verification. As a de-facto standard in LF-VPR, researchers apply epipolar check across images and validate the matched reference by the number of feature inliers dbow-iros ; dbow-tro ; htmap ; AVWs ; BoTW ; fild2 . This method works properly in most scenes but fails if mismatched local features is dominated or pre-defined parametric model is incorrect. So, Yue et al. yue1 ; yue2 and Ma et al. LRO developed new methods to measure geometrical consistency. And Some other works choose to introduce temporal consistency in Eqn. 2 ibuild ; ibow-lcd ; AVWs but this scheme is limited to retrieval in sequential images. They follow a basic assumption that if $I_{q(t_{i})}$ and $I_{r(t_{j})}$ were matched, their adjecent images $I_{q(t_{i}-1)}$ and $I_{r(t_{j}-1)}$ should also be matched.
@@
Sequence Feature-based VPR (SF-VPR): SeqNet seqnet and Delta Descriptor delta-des proposes to fuse temporal information that utilize the changes across image global features as sequence global feature. SeqVLAD SeqVLAD combines NetVLAD layer netvlad and sequence global features that exploits the temporal cues in a sequence. In these works, VPR first performs sequential matching across the image sequences then finds the most matched frame within the matched sequence.
@@
Semantic-based VPR (S-VPR): Gawel et al. xview utilized semantic objects as features, and built a semantic topological graph based on the nearby relationship of semantic objects in images. The graph can be compactly represented by a global feature based on random walk deepwalk and S-VPR is performed in the GF-VPR way.
@@
Summary: VPR algorithms aim to retrieve geographical nearby reference frame. Benefit from robustly representing images/map, current VPR solution can fully and precisely retrieve the frame. However, VPR perform localization by regarding the reference pose as an approximated query pose, which inherently limits the localization accuracy when the current trajectory is very different from the mapping trajectory. Thus, VPR usually serves as a coarse or initial localization step in practical localization system hloc1 ; hloc2 .
3.2 Relative Pose Estimation
@@
Relative Pose Estimation (RPR) methods aims to estimate the relative pose between query image and reference image in the map. In this field, the scene map is also represented by images with pose, i.e., $M=\{\textbf{m}_{t}\}$, where $\textbf{m}_{t}=<\mathcal{I}_{r(t)},\textbf{x}^{\mathcal{G}}_{t}>$. The scene sometimes only contain one reference image $M=\textbf{m}_{r}=<\mathcal{I}_{r},\textbf{x}^{\mathcal{G}}_{r}>$ arnold2022mapfree . When the query image $\mathcal{I}_{q}$ is fed into the system, the RPR algorithm estimate the relative pose of $\mathcal{I}_{q}$ with regard to its (closest) reference image $\mathcal{I}_{r}$:
@@
where *** is the pose transform operator. RPR methods avoid costly pre-scanning and reconstructing scene, enabling economical MRL in some new environment with extremely sparse records, which is commonly seen in Augmented Reality (AR) applications: User A shares a photo with its location in a new scene, any user B can instantly re-localize with regard to user A and perform interactions arnold2022mapfree . Generally, RPR methods can be categorized into geometry-based methods and learning-based methods. This section will review RPR algorithms from such a perspective.
Geometry-based RPR (G-RPR): The G-RPR methods perform feature matching between query image and reference image to obtain pixel-level correspondence, and then heuristically solve relative pose. According to multiple view geometry theory, the relative pose (relative rotation and scaleless translation) between two images with known camera intrinsics can be recovered by decomposing the essential matrix MVG . The essential matrix is classically estimated by matching local features, followed by robust estimation, such as using a 5-point solver 5points inside a RANSAC circulation ransac . Later proposals further improve such a basic formula by learning better features sift ; surf ; orb ; superpoint ; disk ; alike ; aliked ; r2d2 , better matching superglue ; LoFTR ; adalam ; lindenberger2023lightglue and better robust estimators ransac_analysis ; good_correspondence ; MAGSAC ; MAGSAC++ ; NGRANSAC ; ACNe .
@@
In the early G-RPR solutions, scholars utilized scaleless pairwise relative poses between the query image and multiple reference images to triangulate the scaled, metric pose of query image EssNet ; ExReNet . However, in the scene where only one reference image exists, such a scheme fails to recover scale MVG . In arnold2022mapfree , as show in Fig. 6, authors proposed to utilized estimated monocular depth DPT ; planeRCNN to a) provide scale or b)lift 2D-2D correspondences to 2D-3D or c) 3D-3D correspondences, and concluded the second solution back-projecting reference image to 3D space achieves best performance, which is technically converting G-RPR to typical Perspective-n-Point (PnP) problem. And arnold2022mapfree also concludes better feature matching could boost the G-RPR methods.
Learning-based RPR (L-RPR): With the help of DNN model, L-RPR algorithms predict the relative pose from two input images in an E2E way and bypass explicit estimation of 2D correspondences. Some methods first estimate pose up to scale and then recover scale by triangulation CNNSPP ; ExReNet ; Zakaria2017ICCVW , while others aim to directly estimate metric relative pose DistillPose ; RelocNet ; RPNet . Both RelocNet RelocNet and ExReNet ExReNet show great generalization ability when perform RPR on unseen scenarios. Since L-RPR methods do not need to match between query image and reference image, they generally have low requirements of texture in the scene. Thus, L-RPR has been applied in scenarios that are challenging for G-RPR approaches. Cai et al. Cai2021ExtremeRot focused on estimating the relative rotation between two images in extreme cases, including when there is no overlap between the two images. Similarly, the method in DirectionNet estimates scaleless relative pose for pairs of images with very low overlap.
@@
In arnold2022mapfree , authors combined two type of PRP methods and designed a network to predict 3D-3D correspondence between two images, adopting KeypointNet keypointnet . The relative pose is then solved as G-RPR method using Orthogonal Procrustes Procrustes inside a RANSAC loop ransac . The experimental results claimed that this solution outperforms the L-RPR method, indicating that involving multiple view geometry knowledge in neural network may benefit RPR algorithms.
@@
Summary: RPR perform localize by estimating relative pose between query and reference images. Since scale recovery for pairwise RPR is not perfectly solved yet, RPR methods cannot achieve very high-precision localization performance. But, the low requirement of scene map still let RPR attract growing attention in light applications like AR.
4 Visual Landmark Map
As the most popularly applied representation format of scene map, visual landmarks is constructed by Structure from Motion (SfM) colmap or SLAM orbslam3 . Visual landmarks is some informative and representative 3D points that lifted from 2D pixels by 3D reconstruction, and they are associated with corresponding local features in various observed images including 2D key point and high-dimensional descriptor as shown in Fig. 7, i.e., scene map can formatted as =$\mathcal{M}=\{\textbf{m}_{i}\}$, where $\textbf{m}_{i}=<\textbf{p}^{\mathcal{G}}_{i},\{\mathcal{F}^{t}_{i}\}_{t}>$ and $\mathcal{F}^{t}_{i}$ is a set of local features from reference image $\mathcal{I}_{t}$ associated with $\textbf{m}_{i}$, $\textbf{p}^{\mathcal{G}}_{i}$ denotes the 3D location of $\textbf{m}_{i}$ in global frame $\mathcal{G}$. Among the Visual Landmarks-based MRL methods (VL-MRL), Hierarchical Localization (HLoc) hloc1 ; hloc2 is the most famous framework and has been widely used in many applications. Up to now, it still dominates the long term visual localization task [^1] https://www.visuallocalization.net/benchmark/.
@@ 4 Visual Landmark Map
In HLoc framework hloc1 ; hloc2 , scene map $\mathcal{M}$ composed by 3D visual landmarks with descriptors is offline built via SfM reconstruction colmap . During online localization stage, query image $\mathcal{I}_{q}$ is matched with retrieved reference image $\mathcal{I}_{r}$ and the resulting 2D-2D matches are lifted to 2D-3D matches between $\mathcal{I}_{q}$ and scene map based on $\mathcal{M}$, which can be used to solve scaled pose as a typical PnP problem. For clarify, we split VL-MRL algorithms into two step based on their functions (Step 1.A and 1.B are alternative):
@@
    $\begin{split}\mathcal{F}^{q}=\{\textbf{f}^{q}_{i}\}_{i}&=F_{\mbox{feat}}(\mathcal{I}_{q})~{}~{}\mbox{during localization}\\ \mathcal{F}^{r}=\{\textbf{f}^{r}_{i}\}_{i}&=F_{\mbox{feat}}(\mathcal{I}_{r})~{}~{}\mbox{during mapping}\end{split}$ (8)
@@
    $\{\textbf{f}^{q}_{i},\textbf{f}^{r}_{i}\}_{i}=\mbox{match}(\mathcal{F}^{q},\mathcal{F}^{r})$ (9)
    where $i$ is the indices of matched local features.
@@
    or jointly extract and match local feature between $\mathcal{I}_{q}$ and $\mathcal{I}_{r}$:
    $\{\textbf{f}^{q}_{i},\textbf{f}^{r}_{i}\}_{i}=F_{\mbox{fm}}(\mathcal{I}_{q},\mathcal{I}_{r})$ (10)
    where $i$ is the indices of matched local features.
@@
    $\begin{split}\hat{\textbf{x}}^{\mathcal{G}}_{q}&=\mbox{solvePnP}(\{\textbf{f}^{q}_{i},\textbf{f}^{r}_{i}\}_{i}~{}|~{}\mathcal{M})\\ &=\mbox{solvePnP}(\{\textbf{f}^{q}_{i},\textbf{f}^{\mathcal{G}}_{i}\}_{i})\end{split}$ (11)
    where $i$ is the index of matched local features or visual landmarks.
@@
In this section, we review related proposals in the two steps of VL-MRL methods, namely, local feature extraction then matching (Step 1.A), joint local feature extraction and matching (Step 1.B), and pose solver (Step 2). And we also review some proposals aiming at other perspectives in VL-MRL methods.
4.1 Local Feature Extraction-then-Matching

### v3

@@ III-A Visual Place Recognition
Meanwhile, many other attempts are made by scholars to boost the representative ability of global feature. GeM [63] designs a trainable generalized-mean pooling layer that generalizes maximum and average pooling for feature aggregation. TransVPR [64] could aggregate task-relevant features by self-attention mechanism in Transformers [65]. MixVPR [66] incorporates a global relationship between elements in each feature map in cascaded feature mixing modules. AnyLoc [42] leverages an off-the-shelf self-supervised visual foundation model features (such as DINOv2 [67]) with no VPR-specific training or finetuning.
Although global feature can compactly represent images, the retrieval time of GF-VPR methods will linearly increase with the size of the database. And the robustness against viewpoint changes and dynamic occlusions of global features is also inferior than local feature [68].
Local Feature-based VPR (LF-VPR): A local feature algorithm extracts massive key points and corresponding descriptors from images. Then, the extracted local features are utilized to retrieve the best-matched reference frames. In early LF-VPR researches, $\mbox{ImFeat}(\cdot)$ is traditional hand-crafted local features [56, 69, 70]. Then, DNN-based local features are adopted in VPR tasks. Early researches directly apply off-the-shelf DNN model pretrained by other computer vision tasks as local feature extraction module [71, 72, 73], and recent methods leverage localization task-specific local features and achieve improved performances [57, 74]. In dynamic scene, local features located at dynamic objects/regions will confuse VPR methods, so Chen et al. [34] removed inconsistent dynamics by exploiting instance-level semantics. But semantics are hard to accurately judge the actual motion state of the objects, so DSFeat [68] learns to select stable and discriminative local features based on not only semantics but also attentions.
Directly matching between local features and using overall matching similarity is a straightforward way to measure image similarity, but it is computational costly. Local features need to be further processed to measure image similarity, so the design of $P(A,B)$ is the core component of LF-VPR methods. One common strategy is to aggregate extract local features to a compact global feature, e.g., VLAD [55, 75], ASMK [61], and BoW [60], and performance retrieval as GF-VPR methods. As typical solutions. DBoW [76, 77] builds tree-shaped vocabulary to aggregate local features so that an image can be compactly represented by the occurrences of visual words (some general local features) in the vocabulary, following the principle of BoW [60]. And in [78], authors adopted ASMK-based framework[61] to aggregate local features and retrieve candidates. Compared with the BoW [60], ASMK [61] can eliminate the problem of quantization noise and achieve improved precision. In addition to those LF-VPR solution aggregating local features and retrieving like GF-VPR methods, some other methods adopt probabilistic scheme to find matched reference frame in $P(A,B)$. FAB-MAP [79] and FAB-MAP 2.0 [80] train a Chow-Liu tree to learn the co-visibility of features and then estimate the matching probability of candidates. IBuILD [81] tracks local features along video streams to incrementally generate vocabulary, and perform VPR by a likelihood function. iBoW-LCD [82] applies Bayesian framework by measuring the importance of local features with regard to the vocabulary and query image. Then, LiPo-LCD [83] improves by adding line-wise local feature in human-made environments. Tsintotas et al. found reference images by a binomial probability density function where local features are assigned to images as votes [84, 85].
@@
Merely using descriptor information of local features will confuse distinct places with visually similar textures, called perceptual aliasing problem as show in Fig. 3 d). So, many advanced methods exploit geometrical information from key points of local features as a re-ranked matching score or post verification after obtaining candidate reference images by Eqn. 3. Patch-NetVLAD [20] matches patch-level local DNN features as geometrical verification to re-order candidates retrieved by NetVLAD [19]. As a de-facto standard in LF-VPR, researchers apply epipolar check across images and validate the matched reference by the number of feature inliers [76, 77, 86, 84, 85, 74]. This method works properly in most scenes but fails if mismatched local features were dominated or pre-defined parametric model was incorrect. So, topological graph-based post-verification [87, 88] and local relative orientation matching et al. [78] are developed to measure geometrical consistency. Some other works introduce temporal consistency as post-verification [81, 82, 84], but this scheme is limited in image sequence. They follow a basic assumption that if $I_{q(t_{i})}$ and $I_{r(t_{j})}$ were matched, their adjacent images $I_{q(t_{i}-1)}$ and $I_{r(t_{j}-1)}$ should also be matched.
Figure 6: Three G-RPR solution proposed in [21], namely, a) first solve scaleless relative pose by 2D-2D matches then recover scale by 3D-3D matches using estimated depth, b) first lift 2D-2D matches to 2D-3D matches by depth then estimate scaled pose by PnP+RANSAC solver, c) lift 2D-2D matches to 3D-3D matches using depth then estimate scaled pose.
@@
Sequence Feature-based VPR (SF-VPR): SeqNet [89] and Delta Descriptor [90] proposes to measure the changes across image global features as sequence global feature. SeqVLAD [91] combines NetVLAD layer [19] and sequence global features. In these works, VPR first performs sequential matching across the image sequences then finds the most matched frame within the matched sequence. The SF-VPR is technically similar to GF-VPR but applied in image sequences.
@@
Semantic-based VPR (S-VPR): Gawel et al. [39] utilized semantic objects as nodes, and built a semantic topological graph based on the nearby relationship of semantic objects in images. The graph can be compactly represented by a global feature based on random walk [92]. Also, S-VPR is performed as GF-VPR but global feature is calculated by semantics.
@@
Summary: VPR algorithms retrieve geographical nearby reference frame to localize. Benefit from robustly representing images and map, current VPR solution can fully and precisely retrieve the frame. However, VPR achieves localization by regarding the reference pose as an approximated query pose, which inherently limits the localization accuracy when the current trajectory is very different from the mapping trajectory. Thus, VPR usually serves as a coarse or initial localization step in high-precision localization system [23, 24, 6].
III-B Relative Pose Estimation
@@
Relative Pose Estimation (RPR) methods aims to estimate the relative pose between query image and reference image in the map. In this field, the scene map is also represented by images with pose, i.e., $M=\{\textbf{m}_{t}\}$, where $\textbf{m}_{t}=<\mathcal{I}_{r(t)},\textbf{x}^{\mathcal{G}}_{t}>$. The scene sometimes only contain one reference image $M=\textbf{m}_{r}=<\mathcal{I}_{r},\textbf{x}^{\mathcal{G}}_{r}>$ [21]. When the query image $\mathcal{I}_{q}$ is fed into the system, the RPR algorithm estimate the relative pose of $\mathcal{I}_{q}$ with regard to its (closest) reference image $\mathcal{I}_{r}$:
@@
where *** is the pose transform operator. RPR methods avoid costly pre-scanning and reconstructing scene, enabling economical MRL in some new environment with extremely sparse records, which is commonly seen in Augmented Reality (AR) applications: User A shares a photo with its location in a new scene, any user B can instantly re-localize with regard to user A and perform interactions [21]. Generally, RPR methods can be categorized into geometry-based methods and learning-based methods.
Geometry-based RPR (G-RPR): The G-RPR methods perform feature matching between query image and reference image to obtain pixel-level correspondence, and then heuristically solve relative pose. According to Multiple View Geometry (MVG) theory [93], the relationship between the 2D projection of a 3D point on camera coordinates located on two views can be formatted as an essential matrix $E$:
${(K_{1}\textbf{p}_{1})}^{T}E(K_{2}\textbf{p}_{2})=0$ (10)
where $\textbf{p}_{1}$ and $\textbf{p}_{2}$ denotes 2D projection in image coordinate (key point). And the essential matrix actually contains the relative rotation and translation between two views:
$E=\textbf{t}\times\textbf{R}$ (11)
So, G-RPR problem with known camera intrinsic can be solved by first matching local features, then estimating an essential matrix, and finally decomposing the essential matrix to the relative pose (including relative rotation and scaleless translation) between two images. Later proposals further improve such a basic formula by applying better features [56, 70, 69, 57, 94, 95, 96, 97], better matching [98, 99, 100] or better robust estimators [101, 102, 103, 104, 105, 106].
@@
Since the relative pose decomposed from Essential matrix is up-to-scale, so many scholars utilized scaleless pairwise relative poses between the query image and multiple reference images to triangulate the scaled, metric pose of query image in the early G-RPR solutions [107, 108]. However, in the scene where only one reference image exists, such a scheme fails to recover scale. In [21], as show in Fig. 6, authors proposed to utilized estimated monocular depth [109, 110] to a) provide scale or b) lift 2D-2D correspondences to 2D-3D or c) 3D-3D correspondences, and concluded the second solution back-projecting reference image to 3D space and estimating scaled relative pose by 2D-3D matches achieves best performance, which is technically converting G-RPR to typical Perspective-n-Point (PnP) problem and quite similar to MRL with visual landmark map that we will discuss later in Sec. IV. And [21] also experimentally concludes better feature matching could boost the G-RPR methods.
Learning-based RPR (L-RPR): Using DNN model, L-RPR algorithms can predict the relative pose from two input images in an E2E way, eliminating the requirement of explicit local feature matching and pose recovery. Such a scheme decreases the requirement of scene texture and co-visible area between images, so it can work well in some challenging scenarios that G-RPR approaches fails due to disabled module. For example, in [111, 112], authors focused on develop RPR in extreme cases, including when there is low or even no overlap between the query image and reference image. Additionally, DNNs empower L-RPR methods the ability of scale recovery. Despite some methods still estimate pose up-to-scale and recover scales by triangulation [113, 108, 114], there are some proposals directly estimating metric relative pose between query image and reference image [115, 22, 116]. These two strategies can both show pleased generalization ability on unseen scenarios.
@@
In [21], authors combined G-RPR and L-RPR strategies, and designed a network to predict 3D-3D correspondence between two images. The relative pose is then solved as G-RPR method using Orthogonal Procrustes [117] inside a RANdom SAmple Consensus (RANSAC) loop [118]. The experimental results claimed that this solution outperforms the L-RPR method, indicating that involving MVG knowledge in DNN may benefit RPR algorithms.
@@
Summary: RPR perform localization by estimating relative pose between query and reference images. Since scale recovery for pairwise RPR is not perfectly solved yet, RPR methods cannot achieve very high-precision localization performance. But, the low requirement of scene map, and slightly higher localization accuracy than VPR still let RPR attract growing attention in light applications like AR.
@@
IV Visual Landmark Map
As the most popularly utilized representation format of scene map, visual landmark map is generally constructed by Structure from Motion (SfM) [119] or SLAM [120]. Here, visual landmarks are some informative and representative 3D points that lifted from 2D pixels by 3D reconstruction, and they are associated with corresponding local features in various observed images as shown in Fig. 7. Formally, scene map can represented as =$\mathcal{M}=\{\textbf{m}_{i}\}$, where $\textbf{m}_{i}=<\textbf{P}^{\mathcal{G}}_{i},\{\mathcal{F}^{t}_{i}\}_{t}>$ is a visual landmark, $\textbf{P}^{\mathcal{G}}_{i}$ denotes the 3D location of $\textbf{m}_{i}$ in global frame $\mathcal{G}$, $\mathcal{F}^{t}_{i}$ is a set of local features from reference image $\mathcal{I}_{r(t)}$ associated with $\textbf{m}_{i}$. Among the Visual Landmarks-based MRL methods (VL-MRL), Hierarchical Localization (HLoc) [23, 24] is the most famous framework that has been widely used in many applications. Up to now, it still dominates the long-term visual localization task [^1] https://www.visuallocalization.net/benchmark/.
In HLoc framework [23, 24], visual landmark map is offline built via SfM reconstruction [119]. During online localization stage, query image $\mathcal{I}_{q}$ is matched with retrieved reference image $\mathcal{I}_{r}$ and the resulting 2D-2D matches are lifted to 2D-3D matches between $\mathcal{I}_{q}$ and visual landmarks based on $\mathcal{M}$, which can be used to solve scaled pose as a typical PnP problem. For clarify, we split VL-MRL algorithms into two-step pipeline (Step 1.A and 1.B are alternative):
@@ IV Visual Landmark Map
    $\begin{split}\mathcal{F}^{q}=\{\textbf{f}^{q}_{i}\}_{i}&=\mbox{ImFeat}(\mathcal{I}_{q})~{}~{}\mbox{during localization}\\ \mathcal{F}^{r}=\{\textbf{f}^{r}_{i}\}_{i}&=\mbox{ImFeat}(\mathcal{I}_{r})~{}~{}\mbox{during mapping}\end{split}$ (12)
@@
    $\{<\textbf{f}^{q}_{i},\textbf{f}^{r}_{j}>\}=\mbox{Match}(\mathcal{F}^{q},\mathcal{F}^{r})$ (13)
    where $i,j$ are the indices.
@@
    or jointly extract and match local feature between images:
    $\{<\textbf{f}^{q}_{i},\textbf{f}^{r}_{j}>\}=\mbox{FM}(\mathcal{I}_{q},\mathcal{I}_{r})$ (14)
@@
    $\begin{split}\hat{\textbf{x}}^{\mathcal{G}}_{q}&=\mbox{SolvePnP}(\{<\textbf{p}^{q}_{i},\textbf{p}^{r}_{j}>\}~{}|~{}\mathcal{M})\\ &=\mbox{SolvePnP}(\{<\textbf{p}^{q}_{i},\textbf{P}^{\mathcal{G}}_{j}>\})\end{split}$ (15)
@@
In this section, we review related algorithms in these two steps of VL-MRL methods, namely, local feature extraction then matching (Step 1.A), joint local feature extraction and matching (Step 1.B), and pose solver (Step 2). And we also review some proposals aiming at other perspectives in VL-MRL methods.
IV-A Local Feature Extraction-then-Matching